jon and sue are going to elope. Jon brings a ladder to sue window the ladder is 1 foot more than the vertical distance from the ground to the bottom of her window.In order for the ladder to just reach the bottom of her window .Jon should place the bottom of the ladder 9 feet away from the side of the house.How long is the ladder?

Answer by Greenfinch(383)

If distance from ground to window is X, the ladder is (X + 1) and the distance from the wall to the ladder is 9.
So Pythagoras says, X^2 + 9^2 = (X + 1)^2
X^2 + 81 = X^2 + 2X + 1
2X = 80, X=40
So the ladder is 41 feet.
You could also use the fact that Pythagorean triples are such that the small side squared is equal to the other two sides
3^2 = 4 + 5
5^2 = 12 + 13
7^2 = 24 +25
9^2 = 40 + 41